OpenStudy (anonymous):

Which is closest to the area of a circle with a diameter of 19 meters? A. 59.7 square meters B. 29.8 square meters C. 283 square meters D. 890 square meters

OpenStudy (anonymous):

i forgot the formula

OpenStudy (anonymous):

compute \[\pi\times 19^2\]

OpenStudy (anonymous):

go with the decimal approximation here http://www.wolframalpha.com/input/?i=pi*19^2

OpenStudy (anonymous):

1133.54

OpenStudy (mindblast3r):

the answer is A.

OpenStudy (anonymous):

general fromula is \(A=\pi r^2\)

OpenStudy (mindblast3r):

pi x radius^2

OpenStudy (anonymous):

whoa nelly did i ever make a mistake!

OpenStudy (mindblast3r):

radius = half of diameter.

OpenStudy (anonymous):

the diameter is 19 so the radius is \(9.5\) half of it !!

OpenStudy (anonymous):

so a?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=pi*%289.5%29^2

OpenStudy (mindblast3r):

oh wait noooooo i'm wrong satellite is right my bad lol

OpenStudy (anonymous):

nono not A

OpenStudy (anonymous):

it is the usual answer, C

OpenStudy (mindblast3r):

i was going by satellites first formula

OpenStudy (mindblast3r):

yah it's c

OpenStudy (anonymous):

another c realy lol

OpenStudy (anonymous):

just stop asking and put in C for all save a lot of time you will get 80% at least

OpenStudy (mindblast3r):

O.o what are you doing alexis?

OpenStudy (anonymous):

i did it was a joke

OpenStudy (anonymous):

The area of a triangle is 299 square centimeters. The length of the base is 46 centimeters. What is the height of the triangle? A. 6.5 centimeters B. 126.5 centimeters C. 13 centimeters D. 92 centimeters

OpenStudy (mindblast3r):

c

OpenStudy (anonymous):

\[A=\frac{1}{2}bh\] so \[299=\frac{1}{2}\times 46\times h\] or \[299=23h\]

OpenStudy (mindblast3r):

46 x 2 / 2 = 299

OpenStudy (mindblast3r):

B x H / 2 46 x 13 / 2 = 299cm2

OpenStudy (anonymous):

that makes \[h=299\div 23=13\]

OpenStudy (anonymous):

surprise surprise C again

OpenStudy (anonymous):

we did get a D in there one time

OpenStudy (mindblast3r):

lol satellite your more helpful than me your helping her by telling her how to do it, i'm just g iving the answer.

OpenStudy (anonymous):

Which is the area of the figure? A. 192 square inches B. 324 square inches C. 392 square inches D. 480 square inches

OpenStudy (anonymous):

god no we have to break it in two a few parts

OpenStudy (anonymous):

i learned this in 5th grade but that was 3 years ago

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

ok lets compute the big rectangle and then get rid of the small piece the big one has dimensions 24 and 12+4+4=20 so its area is \[24\times 20=480\]

OpenStudy (anonymous):

the part we have to get rid of is \(13\times 12=156\)

OpenStudy (anonymous):

\[480-156=324\]is that a choice?

OpenStudy (anonymous):

yep its b

OpenStudy (anonymous):

oh look, a B heaven's sake

OpenStudy (anonymous):

what the formula for volume?

OpenStudy (anonymous):

it is not volume, it is area area of a rectangle is base times height

OpenStudy (anonymous):

volume of a rectagular prism formula

OpenStudy (anonymous):

\[wlh\] width times length times height

OpenStudy (anonymous):

k i got a 69% :(

OpenStudy (anonymous):

?

OpenStudy (anonymous):

with all that help? i know we did it right together do you have to do it again?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which is a correct way to name the ray whose endpoint is point S and that contains point U?

OpenStudy (anonymous):

question?

OpenStudy (anonymous):

Which is a correct way to name the ray whose endpoint is point S and that contains point U?

OpenStudy (anonymous):

choices?

OpenStudy (anonymous):

should be S something to the right

OpenStudy (anonymous):

ummm a.line st b.line sr c.line ru d.line us

OpenStudy (anonymous):

line st going to the right

OpenStudy (anonymous):

What is the measure of the complement of an angle that measures 16°? A. 164° B. 29° C. 106° D. 74°

OpenStudy (anonymous):

complimentary angles add to 180 180-16=164

OpenStudy (anonymous):

wait hold on

OpenStudy (anonymous):

that was wrong, they add to 90

OpenStudy (anonymous):

90 - 16 = 76 sorry

OpenStudy (anonymous):

not a option

OpenStudy (anonymous):

74

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i must be getting tired, i can't even subtract

OpenStudy (anonymous):

Which are the correct classifications for the triangle? A. acute and isosceles B. obtuse and scalene C. acute and scalene D. obtuse and isosceles

OpenStudy (anonymous):

i thin its c

OpenStudy (anonymous):

hold on it is not C

OpenStudy (anonymous):

k

OpenStudy (anonymous):

two sides are the same, so it is isosceles

OpenStudy (anonymous):

also acute

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is the value of x? A. 103° B. 84° C. 70° D. 77°

OpenStudy (anonymous):

internal angles add up to 180 \[180-42-35=103\] two A's in a row!

OpenStudy (anonymous):

Given ∆STC ~ ∆WBL, what is WL? A. 18.75 B. 37.5 C. 32 D. 24

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

\[\frac{x}{20}=\frac{30}{25}\] \[x=\frac{20\times 30}{25}=24\]

OpenStudy (anonymous):

yeah D

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

@satellite73 would you help me

OpenStudy (anonymous):

Which is closest to the area of a circle with a diameter of 17 meters? A. 53.4 square meters B. 106.8 square meters C. 907.5 square meters D. 226.9 square meters

OpenStudy (anonymous):

\[\pi\times 8.5^2\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=pi*8.5^2

OpenStudy (anonymous):

D again

OpenStudy (anonymous):

d

OpenStudy (anonymous):

The area of a triangle is 323 square centimeters. The length of the base is 38 centimeters. What is the height of the triangle? A. 8.5 centimeters B. 17 centimeters C. 76 centimeters D. 142.5 centimeters

OpenStudy (anonymous):

What is the area of a triangle it is = 1/2 * base * height

OpenStudy (anonymous):

\[323\div 19\]

OpenStudy (anonymous):

b

OpenStudy (anonymous):

i get \(17\)

OpenStudy (anonymous):

Which is the area of the figure? A. 476 square inches B. 436 square inches C. 616 square inches D. 486 square inches